OpenStudy (anonymous):

Help !!:/ Raise the number to the indicated power and express in trigonometric notation [2(cos118degree+ i sin 118 degree)]^4

OpenStudy (anonymous):

cos x+i sin x=e^(i x) (cos x+i sin x)^n=e^(i xn)=cos nx+i sin nx